OpenStudy (anonymous):

d/dy sqrt( x+e^(4y) )

OpenStudy (anonymous):

hi

zepdrix (zepdrix):

hi :)

OpenStudy (anonymous):

i got 1/2(x+e^(4y))*4e^(4y) is this correct?

zepdrix (zepdrix):

This is a partial derivative with respect to y? Or just a normal derivative wrt y? :o

OpenStudy (anonymous):

how will that be different if it's normal or not?

zepdrix (zepdrix):

If it's a partial derivative, x is held constant. If it's a normal derivative, we will be getting a dx/dy term (or x') at some point.

zepdrix (zepdrix):

\[\Large \frac{\partial}{\partial y}\sqrt{x+e^{4y}} \quad=\quad \frac{1}{2\sqrt{x+e^{4y}}}(4e^{4y})\]

zepdrix (zepdrix):

Looks like you've got the right idea, maybe you just made a typo? I don't see the square root in your solution.

zepdrix (zepdrix):

Just so there is no confusion, if it had been a normal derivative with respect to y, we would get:\[\Large \frac{d}{dy}\sqrt{x+e^{4y}} \quad=\quad \frac{1}{2\sqrt{x+e^{4y}}}\left(\frac{dx}{dy}+4e^{4y}\right)\]

OpenStudy (anonymous):

oh so if it's partial then i should consider to take derivative only respect to Y

OpenStudy (anonymous):

but if it's normal then i should consider for whole thing inside of ( )

zepdrix (zepdrix):

ya :) in partial y... x=c. in normal, x is a function, giving us those nasty derivative terms :3

zepdrix (zepdrix):

err x is a variable, blah whatever.

OpenStudy (anonymous):

i've been learning partial for weeks and getting confuse between normal and partial T.T

zepdrix (zepdrix):

oh boy :x

OpenStudy (anonymous):

getting used to do partial and forgetting normal..

zepdrix (zepdrix):

Ya partials seem to make a little more sense XD lol

OpenStudy (anonymous):

haha thank you so much for your help!